OpenStudy (anonymous):

if y =(x*sin^-1x)/(√(1-x^2)) prove that (1-x^2)dy/dx=x+(y/x)

hartnn (hartnn):

first multiply both sides by √(1-x^2) for x sin^-1 x you'll need product rule, know what that is ?

OpenStudy (anonymous):

yeah i know how to do but not getting the answer nah -_- i used division ru;e need help in differenatition of xsin^-1x

OpenStudy (anonymous):

x*1/(1-x^2)*-2x + sin^-1x ???????/

OpenStudy (anonymous):

x*1/√(1-x^2)*-2x + sin^-1x

hartnn (hartnn):

ok, diff. of sin^(-1) x is 1/ (sqrt (1-x^2)) right ? [x sin^(-1)x ]' = x /(sqrt (1-x^2)) + sin^(-1) x only

hartnn (hartnn):

why the -2x ?

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

someone told me to do that

hartnn (hartnn):

oh, but if the function was something like sqrt (1-x^2), the you would get -2x but the function here is just sin^-1 x whose derivative is just 1/ (sqrt (1-x^2))

OpenStudy (anonymous):

oh but still i m not getting teh answer -_-

OpenStudy (anonymous):

i tried so many times

hartnn (hartnn):

no question of chain rule because there's no composite function... show your work, i'll spot the error

OpenStudy (anonymous):

okay brb in 1 min

hartnn (hartnn):

by that time, i'll post my way of doing it, not using quotient rule...

hartnn (hartnn):

y =(x*sin^-1x)/(√(1-x^2)) y √(1-x^2) = x sin^-1 x now differentiating w.r.t x (dy/dx) (√(1-x^2) + y ([-2x]/[2(√(1-x^2)]) = x /(sqrt (1-x^2)) + sin^(-1) x 2 product rules ..... now from the question, y =(x*sin^-1x)/(√(1-x^2)) isolating sin^-1 x, we get sin^-1 x = y √(1-x^2) / x plugging this in, (dy/dx) (√(1-x^2) + y ([-2x]/[2(√(1-x^2)]) = x /(sqrt (1-x^2)) +y √(1-x^2) / x now multiplying both sides by √(1-x^2) , we get (1-x^2)(dy/dx) -xy = x +y(1-x^2)/x --->(1-x^2)(dy/dx) = x +y/x -xy+xy (1-x^2)(dy/dx) = x+y/x thats it!

OpenStudy (anonymous):